OpenStudy (anonymous):

f(-6)=???? f(x)=square root of 5-x

OpenStudy (accessdenied):

\( f(x) = \sqrt{5 - x} \) This is it, correct? So, f(-6) means x = -6 in our function. Have you tried replacing x for -6 in the function?

OpenStudy (anonymous):

yes and it comes as a decimal .. 3.316

OpenStudy (accessdenied):

Hmm, that seems about right. I might leave it with the radical, myself, depending on if the question asks us to round or not. :)

OpenStudy (anonymous):

it wants a radical

OpenStudy (anonymous):

? i dont know how to do it

OpenStudy (accessdenied):

\( f(-6) = \sqrt{5 - (-6)} \) \( = \sqrt{5 + 6} \) \(= \sqrt{11} \) <--- this is the radical, I think I'd leave it there if you're not asked to round \( \approx 3.316\) <--- this is your answer, which is the same.

OpenStudy (accessdenied):

\(\sqrt{11}\) is just a little more "accurate". :)